OpenStudy (anonymous):

The sum of the measures of the interior angles in a polygon is 540 degrees. How many sides does the polygon have? 3 4 5 6

OpenStudy (anonymous):

sup creepy

OpenStudy (anonymous):

ik u

OpenStudy (anonymous):

5

OpenStudy (anonymous):

se the theorem that says that if n is the number of sides of a polygon the sum of the interior angles is (n-2)*180. So we have: (n-2)*180 = 540 180n - 360 = 540 180n = 540 + 360 180n = 900 n = 5

OpenStudy (anonymous):

ok thnx :)